# Puzzle Solving using Reasoning of Large Language Models: A Survey

### v2

@@ 1 Introduction
Figure 2: Taxonomy of Puzzles
@@ 3.1 Prompting Methods
Recent works focus on how different ‘thought structures’ can guide LLMs to the final solution.
@@
Chain topologies, which include Chain-of-Thought (CoT) Wei et al. (2022); Kojima et al. (2022) have been applied to all kinds of puzzles, demonstrating their superiority over simple IO prompts. Self-Refine Madaan et al. (2023) is used for the Game of 24 (rule-based/deterministic), outperforming CoT with a 13% higher success rate Yao et al. (2023). Gu et al. (2023) use Automatic CoT Zhang et al. (2022), Complexity CoT Zhang et al. (2022) and Plan-and-Solve Wang et al. (2023a) in a rule-less detective-style benchmark, with none of the methods clearly outperforming CoT across all tested LLMs. The best results are achieved by Detective Thinking Prompt, a CoT-like method introduced in the same study, which does not exceed the 61.6% accuracy score of the best model, GPT-4. Schuster et al. (2021) exclusively utilized the solutions to programming puzzles that the model had already solved as examples, surpassing alternative approaches.
@@
Tree topologies cover a variety of methods. Self-Consistency (SC) Wang et al. (2022) has been tested on rule-based/deterministic puzzles, such as the 8-puzzle, Game of 24 and Pocket Cube, as well as on rule-less commonsense reasoning puzzles, showcasing a small gain in the first category over CoT Ding et al. (2023); Yao et al. (2023); Mo and Xin (2023) and no clear benefit in the second one Gu et al. (2023). Tree-of-Thought(s) (ToT) Yao et al. (2023); Long (2023) has been exclusively applied to rule-based/deterministic puzzles so far, achieving significantly improved success rates over CoT, with increases ranging from 26% Mo and Xin (2023) to 70% Yao et al. (2023) depending on the puzzle and the depth of the tree, despite the increased LLM invocations Ding et al. (2023). Tree-of-Uncertain-Thought (TouT) Mo and Xin (2023) achieved even better results than ToT on the same challenges, with a 9% higher success rate on the Game of 24 and 3% on mini-crosswords. Finally, Inference-Exclusion-Prompting (IEP) Tong et al. (2023) delivered some of the best results on riddles and commonsense puzzles when combined with CoT, scoring 82% on puzzles–up from 81% with zero-shot CoT–and 79% on riddles, compared to 82% with zero-shot CoT.
@@
Graph topologies entail the following: Graph-of-Thought(s) (GoT) Besta et al. (2023); Lei et al. (2023) and Everything-of-Thought (XoT) Ding et al. (2023) have been used to solve rule-based/deterministic puzzles. While GoT has shown poorer results compared to ToT, with a decrease ranging from 2% to 6% Ding et al. (2023), XoT has been recognized as the most effective method for these puzzles, achieving improvements in results from 53% to 69% compared to ToT, while presenting the fewest LLM invocations among the methods tested, including CoT, SC, ToT, and GoT.
@@
A brief analysis of the aforementioned thought structures is presented in Appendix A.1, while a more detailed one can be found in the work of Besta et al. (2024). Beyond the aforementioned methods, the use of extra information such as hints for riddles and commonsense puzzles, or introductions and summarizations of the puzzles, has also been employed. The inclusion of supplementary details appears to yield positive results, although this is not always the case; for instance, Chinese riddles typically show worse results when hints are used Zhang and Wan (2021).
@@ 3.3 Fine-Tuning
Fine-tuning LLMs emerges as a potent strategy for enhancing their reasoning capabilities, ranging from general logical reasoning to specific puzzle-solving skills. Models such as LoGiPT Feng et al. (2023a) and LogiT5 Luo et al. (2023) demonstrate improved logical reasoning, mimicking human-like problem-solving processes. In the realm of riddles, the study of Lin et al. (2021) illustrates that models like BERT Devlin et al. (2019), RoBERTa Liu et al. (2019) and ALBERT Lan et al. (2019) perform better when trained on both RiddleSense Lin et al. (2021) and CommonsenseQA Talmor et al. (2019) datasets, leveraging commonsense knowledge effectively. Moreover, Zhang and Wan (2021) report that combining fine-tuning on ALBERT-XXL with transfer learning from CommonsenseQA achieved the highest accuracy, noting a 4% improvement over simple fine-tuning. In the domain of rule-based deterministic puzzles, Noever and Burdick (2021) observe suboptimal results when fine-tuning GPT-2 on Sudoku, Rubik’s Cube and Mazes, potentially due to a brief fine-tuning period and limited training examples. Regarding crosswords, various studies Rozner et al. (2021); Efrat et al. (2021) show mixed results, with some fine-tuned LLMs outperforming non-neural baselines and others not, highlighting the inherent challenge of cryptic crosswords for LLMs. Kazemi et al. (2023) demonstrate that fine-tuning LLMs with proofs and CoT under rule-based contexts yields some of the best results. Lastly, the effectiveness of fine-tuning extends to commonsense reasoning Del and Fishel (2022) and programming puzzles Schuster et al. (2021), showcasing its broad applicability across puzzle categories.
@@ 4 Datasets, Benchmarks and Tasks
Exploring diverse datasets, benchmarks, and tasks is crucial for evaluating LLMs in puzzle-solving. This section examines datasets within our puzzle taxonomy, encompassing formats, evaluation metrics, and methodologies. Table 2 provides a detailed summary of datasets utilized across the taxonomy’s categories, organized according to puzzle type. The analysis demonstrates LLMs’ versatility and the impact of techniques discussed in §3.
@@ 4.2.3 Commonsense Reasoning Puzzles
True Detective Del and Fishel (2022) presents detective puzzles in long-form stories, challenging LLMs such as GPT-3.5/4 to draw conclusions. Various methods, including CoT and Golden-CoT are used, revealing difficulties in making final inferences despite all information being available. While Vanilla and CoT approaches perform close to random, Golden CoT shows significantly better accuracy, especially on GPT-4.
@@ 5 Discussion and Future Directions
Applied Methods and Dataset Gaps: Across our puzzle taxonomy (Figure 2), the selection of methods such as few-shot prompting, CoT, introductions and fine-tuning is common across most categories. Rule-based deterministic and rule-less commonsense puzzles show the greatest methodological variety, while riddles are also see diverse approaches. In contrast, rule-based stochastic and rule-less programming puzzles exhibit less variety, likely due to fewer studies in these areas. The lack of benchmarks for stochastic puzzles prompted us to include tasks like card and social deduction games, which share core characteristics with traditional puzzles. This highlights the need for more specialized datasets that adhere closely to defined puzzle structures with missing information elements. Additionally, neuro-symbolic techniques that translate natural language into code remain notably underutilized in puzzle benchmarks, suggesting a potential area for future exploration.
@@ A.1 Prompting Topologies
Automatic CoT (auto-CoT) autonomously generates diverse reasoning chains for various questions Zhang et al. (2022), while the complexity of prompted chains influences accuracy, as more intricate reasoning steps often enhance performance in complex inference tasks Fu et al. (2022). This entails generating diverse reasoning chains and selecting outcomes that showcase deeper reasoning capabilities.
Golden CoT offers ground-truth reasoning chains to address limitations of basic prompting, reducing model hallucination risks Del and Fishel (2022). The Plan-and-Solve (PS) method breaks down tasks into subtasks for more structured solving Wang et al. (2023a), while Self-Question guides models through a four-step process to enhance informal reasoning Gu et al. (2023).
@@
Detective Thinking Prompt Gu et al. (2023) encourages the model to consider and analyze multiple clues or pieces of evidence within a given scenario, sequentially building towards a conclusion, much like solving a mystery. This type of prompting can help the model to handle complex scenarios where synthesizing disparate information correctly is crucial to generating accurate and logical outcomes.
@@
The Everything of Thoughts (XoT) framework integrates Monte Carlo Tree Search (MCTS) with LLMs for enhanced thought generation, showing remarkable performance in complex puzzles Ding et al. (2023). Additionally, Inference-Exclusion Prompting (IEP) employs a combination of forward and backward reasoning to approximate human logic more closely Tong et al. (2023).
@@ A.3 Tables
Table 2 summarizes the curated datasets and tasks associated with each category within our taxonomy of puzzles. A detailed examination reveals a substantial number of datasets for rule-based deterministic puzzles, such as Sudoku and Rubik’s Cube, and a variety of rule-less riddles, indicating a strong research interest and resource availability in these areas. However, there appears to be a scarcity in the collection of rule-based stochastic puzzles and rule-less programming puzzles. This gap points to an opportunity for further research and dataset creation that could provide more diverse challenges for advancing the problem-solving capabilities of Large Language Models. Addressing this gap could lead to a more balanced and comprehensive set of benchmarks that reflect a wider spectrum of puzzle-solving scenarios, potentially catalyzing advancements in LLMs’ abilities to handle uncertainty and complex logic-based problem-solving.
Category Type Datasets
Rule-based Deterministic BoardgameQA Kazemi et al. (2023), Sudoku Noever and Burdick (2021); Long (2023); Ishay et al. (2023), Rubik’s Cube Noever and Burdick (2021); Ding et al. (2023), Maze Noever and Burdick (2021), Crossword Yao et al. (2023); Rozner et al. (2021); Efrat et al. (2021); Kulshreshtha et al. (2022), 8-puzzle Ding et al. (2023), Game of 24 Ding et al. (2023); Yao et al. (2023), Chess Ishay et al. (2023); Feng et al. (2023b)
Stochastic Minesweeper Li et al. (2023), BoardgameQA Kazemi et al. (2023), Card Games Huang et al. (2024); Gupta (2023), Social Deduction Games Wang et al. (2023b); Xu et al. (2023b); Lan et al. (2023)
Rule-less Riddles BrainTeaser Jiang et al. (2023), RiddleSense Lin et al. (2021), BiRdQA Zhang and Wan (2021), CC-Riddle Xu et al. (2022), PUZZLEQA Zhao and Anderson (2023), MARB Tong et al. (2023)
Programming P3 Schuster et al. (2021), Savelka et al. (2023)
Commonsense LatEval Huang et al. (2023b), True Detective Del and Fishel (2022), DetectBench Gu et al. (2023), MARB Tong et al. (2023)
Table 2: Collected Datasets and Tasks for each Category

### v3

@@ 1 Introduction
We keep track of the latest progress in the field of puzzle solving using LLM reasoning in our GitHub https://puzzlellms.github.io/.
{forest}
[…]
Figure 2: A taxonomy of Puzzle Categories with the corresponding Datasets.
@@ 3.1 Prompting Methods
Recent works focus on how different ‘thought structures’ can guide LLMs to the final solution Besta et al. (2024).
@@
Chain topologies, which include Chain-of-Thought (CoT) Wei et al. (2022); Kojima et al. (2022) have been applied to all kinds of puzzles, demonstrating their superiority over simple IO prompts. Self-Refine Madaan et al. (2023) is used for the Game of 24 (rule-based/deterministic), outperforming CoT with a 13% higher success rate Yao et al. (2023). Gu et al. (2023) use several methods in a rule-less detective-style benchmark, including Automatic CoT, which autonomously generates diverse reasoning chains for various questions Zhang et al. (2022); Complexity CoT leverages the complexity of prompted chains, where more intricate reasoning steps often lead to improved performance in complex inference tasks by selecting outcomes that demonstrate deeper reasoning capabilities Fu et al. (2022); and the Plan-and-Solve (PS) method, which uses two prompts for each problem—one to generate the reasoning process and the corresponding answer, and another to extract the final answer from the initial generation Wang et al. (2023a). Despite the varied approaches, none of these methods clearly outperformed CoT across all tested LLMs. The best results are achieved by Detective Thinking Prompt, a CoT-like method introduced in the same study, which does not exceed the 61.6% accuracy score of the best model, GPT-4. The method encourages the model to consider and analyze multiple clues within a given scenario, sequentially building towards a conclusion. This type of prompting can help the model handle complex scenarios where synthesizing disparate information correctly is crucial to generating accurate and logical outcomes. Schuster et al. (2021) exclusively utilized the solutions to programming puzzles that the model had already solved as examples, surpassing alternative approaches.
@@
Tree topologies cover a variety of methods. Self-Consistency (SC) Wang et al. (2022) has been tested on rule-based/deterministic puzzles, such as the 8-puzzle, Game of 24 and Pocket Cube, as well as on rule-less commonsense reasoning puzzles, showcasing a small gain in the first category over CoT Ding et al. (2023); Yao et al. (2023); Mo and Xin (2023) and no clear benefit in the second one Gu et al. (2023). Tree-of-Thought(s) (ToT) Yao et al. (2023); Long (2023) has been exclusively applied to rule-based/deterministic puzzles so far, achieving significantly improved success rates over CoT, with increases ranging from 26% Mo and Xin (2023) to 70% Yao et al. (2023) depending on the puzzle and the depth of the tree, despite the increased LLM invocations Ding et al. (2023). Tree-of-Uncertain-Thought (TouT) Mo and Xin (2023) achieved even better results than ToT on the same challenges, with a 9% higher success rate on the Game of 24 and 3% on mini-crosswords. Finally, Inference-Exclusion-Prompting (IEP) Tong et al. (2023) employs a combination of forward and backward reasoning to approximate human logic and delivered some of the best results on riddles and commonsense puzzles when combined with CoT, scoring 82% on puzzles–up from 81% with zero-shot CoT–and 79% on riddles, compared to 82% with zero-shot CoT.
@@
Graph topologies entail the following: Graph-of-Thought(s) (GoT) Besta et al. (2023); Lei et al. (2023) and Everything-of-Thought (XoT) Ding et al. (2023) have been used to solve rule-based/deterministic puzzles. While GoT has shown poorer results compared to ToT, with a decrease ranging from 2% to 6% Ding et al. (2023), XoT has been recognized as the most effective method for these puzzles. XoT integrates Monte Carlo Tree Search (MCTS) with LLMs for enhanced thought generation, achieving improvements in results from 53% to 69% compared to ToT. Additionally, XoT presents the fewest LLM invocations among the methods tested, including CoT, SC, ToT, and GoT.
@@
A brief analysis of some basic methods not described here is presented in Appendix A.1, while a more detailed analysis of all the methods discussed can be found in the extensive work of Besta et al. (2024). Beyond the aforementioned methods, the use of extra information such as hints for riddles and commonsense puzzles, or introductions and summarizations of the puzzles, has also been employed. The inclusion of supplementary details appears to yield positive results, although this is not always the case; for instance, Chinese riddles typically show worse results when hints are used Zhang and Wan (2021).
@@ 3.3 Fine-Tuning
Fine-tuning LLMs emerges as a potent strategy for enhancing their reasoning capabilities, ranging from general logical reasoning to specific puzzle-solving skills.
Logical Reasoning
LoGiPT Feng et al. (2023a) is a language model fine-tuned to excel in logical reasoning tasks by mimicking the symbolic reasoning of logical solvers. The fine-tuning process involves constructing an instruction-tuning dataset comprising natural language (NL) logical questions paired with symbolic reasoning steps. It is fine-tuned to bypass syntax errors typically encountered in NL to symbolic language parsing, enabling it to directly produce answers. LogiT5 Luo et al. (2023) leverages a multi-task learning approach, incorporating diverse datasets to enhance its reasoning capabilities across different logical domains. The model is fine-tuned on the LOGIGLUE benchmark, which includes various logical reasoning datasets, enabling it to perform better on tasks with limited data by transferring knowledge across tasks.
Rule-based Puzzles
In the domain of rule-based deterministic puzzles, Noever and Burdick (2021) observe suboptimal results when fine-tuning GPT-2 on Sudoku, Rubik’s Cube and Mazes, potentially due to a brief fine-tuning period and limited training examples. Regarding crosswords, various studies Rozner et al. (2021); Efrat et al. (2021) show mixed results, with some fine-tuned LLMs outperforming non-neural baselines and others not, highlighting the inherent challenge of cryptic crosswords for LLMs. Kazemi et al. (2023) demonstrate that fine-tuning LLMs with proofs and CoT under rule-based contexts yields some of the best results.
Rule-less Puzzles
In the realm of riddles, the study of Lin et al. (2021) illustrates that models like BERT Devlin et al. (2019), RoBERTa Liu et al. (2019) and ALBERT Lan et al. (2019) perform better when trained on both RiddleSense Lin et al. (2021) and CommonsenseQA Talmor et al. (2019) datasets, leveraging commonsense knowledge effectively. Moreover, Zhang and Wan (2021) report that combining fine-tuning on ALBERT-XXL with transfer learning from CommonsenseQA achieved the highest accuracy, noting a 4% improvement over simple fine-tuning. Lastly, the effectiveness of fine-tuning extends to commonsense reasoning Del and Fishel (2022) and programming puzzles Schuster et al. (2021), showcasing its broad applicability across puzzle categories.
@@ 4 Datasets, Benchmarks and Tasks
Exploring diverse datasets, benchmarks, and tasks is crucial for evaluating LLMs in puzzle-solving. This section examines datasets within our puzzle taxonomy, encompassing formats, evaluation metrics, and methodologies. Figure 2 provides a detailed summary of datasets utilized across the taxonomy’s categories, organized according to puzzle type. The analysis demonstrates LLMs’ versatility and the impact of techniques discussed in §3.
@@ 4.2.3 Commonsense Reasoning Puzzles
True Detective Del and Fishel (2022) presents detective puzzles in long-form stories, challenging LLMs such as GPT-3.5/4 to draw conclusions. Various methods, including CoT and Golden-CoT, are applied, revealing difficulties in making final inferences despite all necessary information being available. Golden-CoT provides the model with the reasoning behind the correct answer, so the model only needs to understand this reasoning and extract the answer. While Vanilla and CoT approaches perform close to random, Golden-CoT demonstrates significantly better accuracy, particularly with GPT-4. However, even with Golden-CoT, GPT-3.5 achieves a solve rate of only 63%, whereas GPT-4 matches human solver results (without access to the reasoning behind the answer).
@@ 5 Discussion and Future Directions
Applied Methods and Dataset Gaps: Across our puzzle taxonomy, methods such as few-shot prompting, CoT, use of introductions and fine-tuning are commonly employed across most categories. Rule-based deterministic and rule-less commonsense puzzles show the greatest methodological variety, while riddles are also see diverse approaches. In contrast, rule-based stochastic and rule-less programming puzzles exhibit less variety, likely due to fewer studies in these areas. Figure 2 reveals that a substantial number of datasets are available for rule-based deterministic puzzles, such as Sudoku and Rubik’s Cube, as well as a variety of rule-less riddles. This indicates a strong research interest and resource availability in these domains. However, there appears to be a scarcity of datasets for rule-based stochastic puzzles and rule-less programming puzzles. This gap highlights an opportunity for further research and dataset creation to introduce more diverse challenges for advancing the problem-solving capabilities of LLMs. The lack of benchmarks for stochastic puzzles led us to include tasks like card and social deduction games, which share core characteristics with traditional puzzles involving incomplete information. Additionally, neuro-symbolic techniques that translate natural language into code remain notably underutilized in puzzle benchmarks, suggesting a potential area for future exploration.
Comparatively, prompting methods like CoT and ToT enhance complex reasoning abilities without altering the underlying model parameters, yet their effectiveness on smaller LLMs requires further exploration. Moreover, methods such as ToT or GoT necessitate a greater number of model invocations compared to CoT and XoT, which could impact their efficiency and scalability Ding et al. (2023). Fine-tuning, while fundamentally enhances reasoning by altering model parameters, is constrained by its specificity to particular tasks. For instance, models fine-tuned on CSQA have been observed to perform worst on the RiddleSense dataset than models fine-tuned directly on RiddleSense Lin et al. (2021), and a model fine-tuned on Poker games may not perform well on tasks like solving a Rubik’s cube. Finally, puzzle translation methods primarily test the model’s ability to interpret and rephrase problems rather than directly solve them, which tends to have a more limited impact on complex reasoning compared to the other two categories of methods.